An inductance of 1 henry and a resistance of 1 ohm are connected in series with a constant e.m.f. of E volts. If 10 A after 5 seconds, find E

1
Expert's answer
2022-03-07T14:51:04-0500
"I(t)=\frac{\cal E}{R}\left(1-e^{-Rt/L}\right)"

"{\cal E}=\frac{I(t)R}{\left(1-e^{-Rt/L}\right)}"

"{\cal E}=\frac{10*1}{\left(1-e^{-1*5/1}\right)}=10.07\:\rm V"